1 kg of steam at 7 bar, entropy 6.5 kJ/kg K, is heated reversibly at constant pressure until the temperature is 250℃. Calculate the heat supplied.


1
Expert's answer
2022-03-30T13:04:02-0400

"S_1=S_f+x*S_{fg}"

6.5=1.99190+x*4.71529

x=0.95606

"H_1=H_f+x*H_{fg}"

696.49+0.95606*2065.75=2671.47 (KJ/kg)\

"Q=H_2-H_1"

2953.51-2671.47=282.04 (KJ/kg)